DETECTION AND CHARACTERIZATION OF SCANDIUM IMIDE, ScNH

\underline{T C. STEIMLE}, J. XIN, A. J. MARR and S. A. BEATON, Dept. of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287-1604.

Supersonic molecular beams of scandium imide, ScNH, and deuterated imide, ScND, were produced using a laser ablation/reaction source and interrogated by laser induced fluorescence and dispersed fluorescence spectroscopy. The fine and magnetic hyperfine interactions and Stark effect in the (0,0) A^{2}Π_{r} - X^{2}Σ^{+} transition of ScNH (origin = 15023.033 cm^{−1}) were analyzed. Ground state vibrational spacings for ScNH and ScND were measured. The determined permanent electric dipole moments are μ( A^{2}Π_{r}) = 4.08(7)D and μ( X^{2}Σ^{+}) = 2.28(15)D. The Sc (I = \frac{7}{2} ) magnetic hyperfine structure was analyzed from which a plausible molecular orbital description was made. A comparision with experimental observations^{a}^{,}^{b}and theoretical predictions^{c} for YNH has been made.

aB. Simard, W.J. Balfour, M. Vasseur, and P.A. Hackett, J. Chem. Phys. 93, 4481,(1990).

bibid Presentation TC13 ”45th Ohio State Symposium on Molecular Spectroscopy”, June 11-15, 1990.

cK.K. Das and K. Balsubramanian,, J. Chem. Phys. 93, 6671,(1990).